OpenStudy (anonymous):

Which is less than zero point eight five? nine-tenths eighteen-twentieths twenty-two twenty-fifths four-fifths

OpenStudy (anonymous):

0.85.... 9/10 = .9 18/20=.9 24/25=.98 4/5=.8 so 4/5

OpenStudy (pixiedust1):

Let's translate these into decimals. How would you do that?